OpenStudy (anonymous):

First timer please help- A company plans to sell embroidered hats for $15 each. The company’s financial planner estimates that the cost, y, of manufacturing the hats is a quadratic function with a y-intercept of 7,920 and a vertex of (150, 9,000). What is the minimum number of hats the company must sell to make a profit? 151 401 529 601

OpenStudy (anonymous):

I chose 529 but it was wrong. I need to know where I went wrong

OpenStudy (amistre64):

you need to construct the quadratic from the given information ...

OpenStudy (amistre64):

to find the break even point, sales have to equal costs

OpenStudy (amistre64):

15x = cost function

OpenStudy (amistre64):

what did you get for your cost function?

OpenStudy (anonymous):

I can't remember my teacher took my work. but she told me that one was wrong..

OpenStudy (anonymous):

I'm just learning. Im the youngest in the class. should be in 8

OpenStudy (amistre64):

well, you have all the information here. how would you construct it? you are given a vertex and a y intercept.

OpenStudy (amistre64):

given a vertex, (n,m) we can construct a vertex form as: y = a(x-m)^2 + n the y intercept is given so that we can fine tune it: for some y intercept (0,k) k = a(0-m)^2 + n k-n = a m^2 (k-n)/m^2 = a

OpenStudy (anonymous):

y=15x and y=0.001(150)^2+9000

OpenStudy (anonymous):

am I right

OpenStudy (amistre64):

my internet is not playing nice :/

OpenStudy (amistre64):

7,920 and a vertex of (150, 9,000). y = a(x-150)^2 + 1000 a = (7920-9000)/(150^2) y = (7920-9000)(x/150-1)^2 + 1000 the wolf tells me that this is a good assessment of it y = -6x^2/125 + 72x/5 -80 therefore our setup is: 15x = -6x^2/125 + 72x/5 -80 0 = -6x^2/125 + 72x/5 -15x -80

OpenStudy (amistre64):

gonna have to rechk my typing :/

OpenStudy (anonymous):

how you get 125

OpenStudy (amistre64):

well for starters, i mistyped 1000 instead of 9000 :)

OpenStudy (amistre64):

expanding out these big numbers is a pain

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

y = (7920-9000)(x/150-1)^2 + 9000 y = -1080(x/150-1)^2 + 9000 y = -1080(x^2/150^2 -2x/150 +1) + 9000 y = -1080(x^2/150^2 -x/125 +1) + 9000 y = -1080x^2/150^2 +1080x/125 -1080 + 9000 y = -1080x^2/150^2 +1080x/125 + 7920 thats what im getting without "simplifying"

OpenStudy (amistre64):

setting that equal to 15x ... then finding the roots :)

OpenStudy (amistre64):

you would find that 400 is one root, so to breakeven they would have to sell 400, to make a profit they would have to sell 401

OpenStudy (anonymous):

ok thank you I think I understand now. Thank you Amistre64, gave you a medal

OpenStudy (amistre64):

youre welcome, good luck :)

OpenStudy (anonymous):

I have another one teacher said was wrong. can you help

OpenStudy (anonymous):

she is pointing to kids who were wrong. embarrassing